answer all questions

2.0B paint nnarcki calmcda. (Dano raund intermediate calculationa. Raund your Tinal anawers to 2 decimal placea.) 3.12 .14
Answer #1

Year

Cash flow

Present value factor

Present value

1

$3.18

0.9009

$2.86

2

$3.66

0.8116

$2.97

3

$4.14

0.7312

$3.03

3

$68.00

0.7312

$49.72

  Total

$58.58

Present value factor = 1 / (1 + r) ^ n

Where,

r = 11% or 0.11

n = Years 1 to 3

Present value of dividend of year 1 = $3.18 × 1 / 1.11= $2.86

Present value of dividend of year 2 = $3.66× 1 / 1.11^2= $2.97

Present value of dividend of year 3 = $4.14 × 1 / 1.11^3= $3.03

Present value of stock price of $68= $68 × 1 / 1.11^3= $49.72